OpenStudy (anonymous):

What is the solution of log3x - 725 = 2? I'm pretty sure i have to convert it to exponential form but I don't know how

OpenStudy (nikato):

\[\log _{3}x - 175 =2 \]

OpenStudy (nikato):

or \[\log 3x - 725=2\]

OpenStudy (anonymous):

neither. it didnt copy right its actually: \[\log_{3x-7} 25=2\]

OpenStudy (nikato):

okay. you can just rewrite it as exponents.  so can you rewrite your equation in expoenents now?